# Minquan Cheng, Huimei Wei, Kai Wan, and Giuseppe Caire

### v2

@@ V-A1 Comparison with the MN scheme in [2]
$\frac{q^{n}-2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}+1}{2^{n}}\approx\frac{q^{n}-(q-1)^{n}}{2^{n}}$
which decreases exponentially with $2$ at a rate of $n$.
@@ V-A2 Comparison with the WCLC scheme in [19]
Compared to the WCLC scheme, our subpacketization increases by a factor of $q$ times while reducing the transmission load by a factor of $2^{n}$.
@@ V-A3 Comparison with the WCWL scheme in [14]
    $\frac{F_{\text{WCWL}}}{F}=\frac{q^{n}}{q^{n}}=1,\ \ \ \ \frac{R_{\text{WCWL}}}{R}=\frac{\frac{2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}(2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}+1)}{2q^{n}}}{\lfloor\frac{q-1}{2}\rfloor^{n}}=\frac{2^{n}(2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}+1)}{2q^{n}}>\frac{2^{n}(q-3)^{n}}{2q^{n}}=\frac{1}{2}\left(2-\frac{6}{q}\right)^{n}.$
@@
    $\frac{R_{\text{WCWL}}}{R}$ $=\frac{\frac{2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}}{2\left\lfloor\frac{q^{n}}{2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}+1}\right\rfloor+1}}{\lfloor\frac{q-1}{2}\rfloor^{n}}=\frac{2^{n}}{2\left\lfloor\frac{q^{n}}{2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}+1}\right\rfloor+1}>\frac{2^{n}(q-3)^{n}}{2q^{n}+(q-1)^{n}+1}>\frac{2^{n}(q-3)^{n}}{3q^{n}}=\frac{1}{3}\cdot\left(2-\frac{6}{q}\right)^{n}.$
@@ V-B Numerical comparisons
In this subsection, we will present numerical comparisons between our scheme introduced in Theorem 3 and the existing schemes that employ linear subpacketization [28, 14, 20, 16, 27], polynomial subpacketization [26, 24], and exponential subpacketization [2, 19]. These schemes are summarized in Table I. It is worth noting that the ASK scheme [27] can be included as a special case of our scheme in Corollary 2 which will be introduced in Subsection VI-A. In addition, the XXGL scheme [28] has large memory ratio which approximates $1$ so in this subsection we do not need to consider the XXGL scheme.
@@ V-B1 Comparison with the linear subpacketization schemes in [20, 16, 14]
In Table II, we present some numerical comparisons with the schemes in [20, 16]. Compared to the ZCW scheme, our scheme in Theorem 3 has more users, a close memory ratio, a similar subpacketization, and a smaller load. Compared to the AST scheme, our scheme has a smaller memory ratio, a lower transmission load, and a smaller subpacketization while having more users.
@@
Now, let us compare our scheme in Theorem 3 with the WCWL scheme in [14]. When $K=85$, we show our scheme (represented by the red line) and the WCWL scheme (represented by the green line) in Figure 3 and Figure 3. We can see that our scheme has a smaller or equal subpacketization, and a lower transmission load for the memory ratio is less than $0.9$. When the memory ratio is close to $1$, our scheme has a slightly higher transmission load while having a smaller or equal subpacketization.
Finally, we would like to point out that the advantages of our scheme in terms of the subpacketization and transmission load becomes more obvious as the number of users $K$ increases. For instance, when $K=729$, we show our scheme and the WCWL scheme in Figure 5 and Figure 5. We can observe that our scheme has a smaller or equal subpacketization and a smaller transmission load.
@@ V-B2 Comparison with the polynomial subpacketization schemes in [26, 24]
Since the user numbers of the schemes in [26, 24] are highly specialized integers, including combination numbers, powers, products of combination numbers, and powers, it is difficult for us to directly plot a graph for given values of $K$. Instead, we compare our scheme with the CKSM scheme and the YTCC scheme by choosing some specific user numbers and memory ratios listed in Table III. We can see that compared to the YTCC scheme, our scheme has a similar memory ratio, a slightly larger transmission load, and more users, while having a much smaller subpacketization. Compared to the CKSM scheme, our scheme has a lower transmission load and a much smaller subpacketization while having a close number of users and a close memory ratio.
@@ V-B3 Comparison with the exponential subpacketization schemes in [2, 19]
Let us first directly compare our scheme in Theorem 3 with the MN scheme and the WCLC scheme, each of which has an exponential subpacketization level with the number of users. When $K=85$, we have our scheme (represented by the red line), the MN scheme (represented by the blue line), and the WCLC scheme (represented by the green line) listed in Figure 7 and Figure 7 by Table I and Theorem 3.
Compared to the MN scheme, we can observe that our subpacketization (the red line) is much smaller than the MN scheme (the blue line), while our transmission load is little larger than the MN scheme when $K=85$. Compared to the WCLC scheme, our scheme (the red line) has a much smaller subpacketization than the WCLC scheme (the green line) while our transmission load is close to or equal to that of the WCLC scheme when $K=85$.
@@
Using the grouping method in Lemma 2 on the schemes in [2, 19], through an exhaustive search on $K_{1}$ given $K=85$, we find that we cannot obtain a scheme with both a lower load and a smaller subpacketization.
Finally we should point out that the advantages on the subpacketization and transmission of our scheme are more obvious as $K$ increases. For instance, when $K=729$ we compare our scheme with the MN scheme and the WCLC scheme in Figure 9 and Figure 9. Note that the ratio of the subpacketization between our scheme and WCLC scheme significantly reduces (see Figure 7 and Figure 9), and our transmission load is more close to that of the WCLC (see Figure 7 and Figure 9).
@@ VI Relationship between NHSDP and the classic combinatorial Designs
In this section, we will show that there is a close relationship between NHSDP and several classical combinatorial designs, such as cyclic difference packing [31], non-three-term arithmetic progressions [32], and perfect hash families [40]. As a result, some new NHSDP can be obtained by the existing cyclic difference packings and some new classes of non-three-term arithmetic progression sets and perfect hash families are obtained by Theorem 3.
@@ VII Conclusion
In this paper, based on the Latin square we introduced a new combinatorial structure called non-half-sum disjoint packing (NHSDP) to construct PDA where the subpacketization is linear with $K$. According to theoretical and numerical comparisons, the proposed scheme achieves lower load than the existing schemes with linear subpacketization; it achieves lower load in some cases than some existing schemes with polynomial subpacketization; it has close loads as some existing schemes with exponential subpacketization in some cases. Furthermore, NHSDP has a close relationship with the classical combinatorial structures such as cyclic difference packing, non-three-term arithmetic progressions and perfect hash family (PHP) and so on. By constructing NHSDPs, we can obtain some new CDPs, NTAPs and PHPs with large cardinalities or columns.
Finally, in this paper we only focus on studying the $(v,g,b)$ NHSDP with $g=2^{n}$ for any positive integer $n$. So it is meaningful to study the $(v,g,b)$ NHSDP for any value of $g$.
@@ Appendix A Proof of Theorem 1
When $d$ runs all the element of $\mathcal{D}_{i}$there exactly $g$ unique solutions. This means that the integer pair $(c,i)$ occurs in $\mathbf{P}$ exactly $g$ times and there are exactly $bv$ different integer pairs. Then the proof is completed.
@@ Appendix B Proof of Lemma 3
Since $v\geq 2\phi(m_{1},m_{2},\ldots,m_{n})+1=2\sum_{i=1}^{n}f(i)+1$, both $\sum_{i=1}^{n}a_{i}x_{i}$ and $\sum_{i=1}^{n}a^{\prime}_{i}x_{i}$ are less tha $\frac{v-1}{2}$. In addition, from (18) for each $i\in[n-1]$ we have $x_{i+1}>2\sum_{j=1}^{i}f(j)$ which implies that

### v3

@@ V-A1 Comparison with the MN scheme in [2]
$\frac{q^{n}-2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}+1}{2^{n}}\approx\frac{q^{n}-(q-1)^{n}}{2^{n}}.$
@@ V-A2 Comparison with the WCLC scheme in [19]
Compared to the WCLC scheme, our subpacketization increases by a factor of $q$ while reducing the transmission load by a factor of $2^{n}$.
@@ V-A3 Comparison with the WCWL scheme in [14]
    $\frac{F_{\text{WCWL}}}{F}$ $=\frac{q^{n}}{q^{n}}=1,$
    $\frac{R_{\text{WCWL}}}{R}$ $=\frac{\frac{2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}(2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}+1)}{2q^{n}}}{\lfloor\frac{q-1}{2}\rfloor^{n}}=\frac{2^{n}(2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}+1)}{2q^{n}}>\frac{2^{n}(q-3)^{n}}{2q^{n}}=\frac{1}{2}\left(2-\frac{6}{q}\right)^{n}.$
@@
    $\frac{R_{\text{WCWL}}}{R}$ $=\frac{\frac{2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}}{2\left\lfloor\frac{q^{n}}{2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}+1}\right\rfloor+1}}{\lfloor\frac{q-1}{2}\rfloor^{n}}=\frac{2^{n}}{2\left\lfloor\frac{q^{n}}{2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}+1}\right\rfloor+1}$
    $>\frac{2^{n}(q-3)^{n}}{2q^{n}+(q-1)^{n}+1}$
    $>\frac{2^{n}(q-3)^{n}}{3q^{n}}$
    $=\frac{1}{3}\cdot\left(2-\frac{6}{q}\right)^{n}.$
@@ V-B Numerical comparisons
In this subsection, we will present numerical comparisons between our scheme in Theorem 3 and the existing schemes with linear subpacketization [28, 14, 20, 16, 27], polynomial subpacketization [26, 24], and exponential subpacketization [2, 19]. These existing schemes are summarized in Table I. It is worth noting that the ASK scheme [27] can be included as a special case of our scheme in Corollary 2 which will be introduced in Subsection VI-A. In addition, the XXGL scheme [28] has large memory ratio which approximates $1$ so in this subsection we do not need to consider the XXGL scheme.
@@ V-B1 Comparison with the linear subpacketization schemes in [20, 16, 14]
Since these schemes have the special integers, including combination numbers, powers, products of combination numbers, and powers, it is difficult for us to directly plot a graph for given values of $K$. So we present some numerical comparisons between our scheme in Theorem 3 and the schemes in [20, 16] in Table II. By Table  II, our scheme in Theorem 3 has more users, a close memory ratio, a similar subpacketization, and a smaller load than that of the ZCW scheme; our scheme has a smaller memory ratio, a lower transmission load, and a smaller subpacketization than that of the AST scheme, while having more users.
@@
Now, let us compare our scheme in Theorem 3 with the WCWL scheme in [14]. When $K=85$, we have our scheme (the red line) and the WCWL scheme (the green line) listed in Figure 3 and Figure 3. We can see that our scheme has a smaller or equal subpacketization, and a lower transmission load than that of the WCWL scheme when the memory ratio is less than or equal to $0.8$. When the memory ratio is close to $0.9$, our scheme has a slightly higher transmission load while having a smaller or equal subpacketization.
Finally, we would like to point out that the advantages of our scheme in terms of the subpacketization and transmission load become more obvious as the number of users $K$ increases. For instance, when $K=729$, we have our scheme and the WCWL scheme listed in Figure 5 and Figure 5. We observe that our scheme has a smaller or equal subpacketization and a much smaller transmission load.
@@ V-B2 Comparison with the polynomial subpacketization schemes in [26, 24]
Since the user numbers of the schemes in [26, 24] are the special integers, including combination numbers, powers, products of combination numbers, and powers, it is difficult for us to directly plot a graph for given values of $K$. Instead, we compare our scheme with the CKSM scheme and the YTCC scheme by choosing some specific user numbers and memory ratios listed in Table III. We can see that compared to the YTCC scheme, our scheme has a similar memory ratio, a slightly larger transmission load, and more users, while having a much smaller subpacketization; compared to the CKSM scheme, our scheme has a lower transmission load and a much smaller subpacketization while having a close number of users and a close memory ratio.
@@ V-B3 Comparison with the exponential subpacketization schemes in [2, 19]
Let us first directly compare our scheme in Theorem 3 with the MN scheme and the WCLC scheme, each of which has an exponential subpacketization level with the number of users. When $K=85$, we have our scheme (the red line), the MN scheme (the blue line), and the WCLC scheme (the green line) listed in Figure 7 and Figure 7 by Table I and Theorem 3.
Compared to the MN scheme, we can observe that our scheme (the red line) has a much smaller subpacketization than the MN scheme (the blue line), while our transmission load is little larger than the MN scheme when $K=85$. Compared to the WCLC scheme, our scheme (the red line) has a much smaller subpacketization than the WCLC scheme (the green line) while our transmission load is close to or equal to that of the WCLC scheme when $K=85$.
@@
Using the grouping method in Lemma 2 on the schemes in [2, 19], through an exhaustive search on $K_{1}$ given $K=85$, we find that we cannot obtain a scheme with both a lower load and a smaller subpacketization. Finally we should point out that the advantages on the subpacketization and transmission of our scheme are more obvious as $K$ increases. For instance, when $K=729$ we compare our scheme with the MN scheme and the WCLC scheme in Fig. 9 and Fig. 9. Note that the ratio of the subpacketization between our scheme and the WCLC scheme significantly reduces (see Fig. 7 and Fig. 9), and our transmission load is more close to that of the WCLC (see Fig. 7 and Fig. 9).
@@ VI Relationship between NHSDP and the classic combinatorial Designs
In this section, we will show that there is a close relationship between NHSDP and several classical combinatorial designs, such as cyclic difference packing [31], non-three-term arithmetic progressions [32], and perfect hash families [40]. As a result, some new NHSDPs can be obtained by the existing cyclic difference packings, and some new classes of non-three-term arithmetic progression sets and some new perfect hash families are obtained by the NHSDPs in Theorem 3.
@@ VII Conclusion
Motivated by the quest to construct coded caching schemes with linear subpacketization and small load, in this work we have introduced a new combinatorial structure called non-half-sum disjoint packing (NHSDP). This structure is used to construct PDAs with linear subpacketization in the number of users K. According to theoretical and numerical comparisons, the proposed scheme achieves lower load than the existing schemes with linear subpacketization; it achieves lower load in some cases than some existing schemes with polynomial subpacketization; it has close loads as some existing schemes with exponential subpacketization in some cases. Furthermore, NHSDP has a close relationship with other classical combinatorial structures such as cyclic difference packing, non-three-term arithmetic progressions, and perfect hash family (PHP). By constructing NHSDPs, we can obtain some new CDPs, NTAPs and PHPs with large cardinalities or columns.
In this paper, we have focused only on $(v,g,b)$ NHSDPs with $g=2^{n}$ for positive integers $n$. In general, we point out as a promising future work to consider also more general NHSDPs with different choice of $g$.
@@ Appendix A Proof of Theorem 1
When $d$ runs all the element of $\mathcal{D}_{i}$, there exactly $g$ unique solutions. This means that the integer pair $(c,i)$ occurs in $\mathbf{P}$ exactly $g$ times and there are exactly $bv$ different integer pairs. Then the proof is completed.
@@ Appendix B Proof of Lemma 3
Since $v\geq 2\phi(m_{1},m_{2},\ldots,m_{n})+1=2\sum_{i=1}^{n}f(i)+1$, both $\sum_{i=1}^{n}a_{i}x_{i}$ and $\sum_{i=1}^{n}a^{\prime}_{i}x_{i}$ are less than $\frac{v-1}{2}$. In addition, from (18) for each $i\in[n-1]$ we have $x_{i+1}>2\sum_{j=1}^{i}f(j)$ which implies that